Four metal rods of lengths 78cm, 104cm, 117cm and 169cm are to be cut into parts of equal length. Each part must be as long as possible. What is the maximum number of pieces that can be cut?
Answer by josgarithmetic(39618)

Try factoring each of the numbers. You may be interested in greatest common divisor.


78=2*3*13;
104=2*2*2*13
117=9*13
169=13*13